# Neural-Quantum-States Impurity Solver for Quantum Embedding Problems

### v1

@@ II.3 The gGA Algorithmic Structure
The gGA theoretical framework is rooted in the variational principle and, similarly to DMFT [7, 8, 9, 10, 11, 12], in the idea of the infinite-dimension limit [65]. Specifically, it is based on a variational wavefunction represented as $|\Psi_{G}\rangle=\hat{P}_{G}|\Psi_{0}\rangle$, where $\hat{P}_{G}$ is an operator acting on a reference single-particle wavefunction (Slater determinant) $|\Psi_{0}\rangle$. The wavefunction $|\Psi_{0}\rangle$ is defined in an auxiliary Hilbert space, where the number of fermionic modes is the product of the number of physical Hilbert space’s fermionic modes and a tunable factor $B$. This factor $B$ governs the size of the variational space and the accuracy of the method. The operator $\hat{P}_{G}$ maps states from the auxiliary space into the physical space, modifying the weights of the local electronic configurations of all fragments $i$. Both $|\Psi_{0}\rangle$ and $\hat{P}_{G}$ are optimised to minimise the variational energy, evaluated neglecting all contributions that vanish in the limit of infinite-dimensional systems (as in DMFT). Therefore, like DMFT, the ghost-GA can accurately capture dynamical correlations, but not spatial correlations.
Here, we provide a brief overview of its algorithmic structure, with further technical details summarised in the appendix. We refer interested readers to the original papers cited above. For simplicity, we specialise the presentation focusing on the single-orbital Anderson lattice model:
$\hat{H}$ $=\sum_{\langle i,j\rangle}\sum_{\sigma}(t_{ij}+\delta_{ij}\epsilon_{p})\,\hat{p}^{\dagger}_{i\sigma}\hat{p}_{j\sigma}+\sum_{i}\hat{H}_{\text{loc}}^{i}\left[d_{i\sigma},d_{i\sigma}^{\dagger}\right]$ (5)
$+V\sum_{i\sigma}\left(\hat{p}^{\dagger}_{i\sigma}\hat{d}_{i\sigma}+\text{H.c.}\right)+\mu\sum_{i}\hat{N}_{i}\,$,
where:
$\hat{H}_{\text{loc}}^{i}\left[d_{i\sigma},d_{i\sigma}^{\dagger}\right]=\frac{U}{2}\left(\hat{n}_{di}-1\right)^{2}\,.$ (6)
Here the Fermionic operators $\{\hat{p}_{i\sigma}^{\dagger},\hat{p}_{i\sigma}\}$ represent the uncorrelated degrees of freedom, while $\{\hat{d}_{i\sigma}^{\dagger},\hat{d}_{i\sigma}\}$ represent the interacting orbitals; $i,j$ are site labels, $\sigma$ is the spin, $\mu$ is the chemical potential, and $\langle i,j\rangle$ indicating a sum over the first nearest neighbours. $\hat{n}_{di}=\sum_{\sigma}\hat{d}_{i\sigma}^{\dagger}\hat{d}_{i\sigma}$, $\hat{n}_{pi}=\sum_{\sigma}\hat{p}_{i\sigma}^{\dagger}\hat{p}_{i\sigma}$ and $\hat{N}_{i}=\hat{n}_{di}+\hat{n}_{pi}$. We focus on the paramagnetic phase and adopt uniform hopping terms $t_{ij}$, resulting in a semi-circular density of states characteristic of the infinite Bethe lattice, with energy measured in units of the half-bandwidth, consistent with previous studies [65].
The algorithmic structure of gGA, resulting from a convenient reformulation of the energy minimisation problem for the multiconfigurational wavefunction $|\Psi_{G}\rangle$ [28, 20, 21], closely mirrors that of DMFT, featuring a self-consistent iterative procedure nested within the solution of an embedding Hamiltonian (EH), which serves as the impurity model:
$\hat{H}^{\text{emb}}[\mathcal{D},\Lambda^{c}]$ $=\sum_{a=1}^{B}\sum_{\sigma=\uparrow\downarrow}\left([\mathcal{D}_{i}]_{a}d_{i\sigma}^{\dagger}b_{ia\sigma}+\text{H.c.}\right)$ (7)
$+\sum_{a,b=1}^{B}\sum_{\sigma=\uparrow\downarrow}[\Lambda_{i}^{c}]_{ab}b_{ib\sigma}b_{ia\sigma}^{\dagger}$
$+\sum_{\sigma=\uparrow\downarrow}\hat{H}_{\text{loc}}^{i}\left[d_{i\sigma},d_{i\sigma}^{\dagger}\right]\,.$
The variational optimisation of the parameters $\mathcal{D}_{i}$ and $\Lambda_{i}^{c}$ ensure that the feedback between local electronic correlations and the global electron density is consistently accounted for, describing how each atomic fragment interacts with its environment.
Once convergence is achieved, observables such as the ground-state total energy of Eq. (5) and the spectral function can be explicitly computed from the optimised variational parameters.
The ghost-GA algorithmic structure outlined above has two key computational advantages: (I) In gGA, it is sufficient to solve an EH with a relatively small bath ($B=3$) to obtain the level of accuracy that DMFT achieves with infinite bath degrees of freedom. (II) While in DMFT it is necessary to compute local Green’s function of the fragment $i$, the ghost-GA requires only the iterative calculation of the entries of the following ground-state single-particle density-matrix entries of the EH: $\langle\Phi|b^{\dagger}_{ia\sigma}b_{ib\sigma}|\Phi\rangle$ and $\langle\Phi|d^{\dagger}_{i\sigma}b_{ib\sigma}|\Phi\rangle$.
However, solving the EH remains the primary computational bottleneck. In the following section, we benchmark our proposed neural quantum state framework against exact diagonalisation, demonstrating its accuracy and supporting its feasibility in future gGA applications to multi-orbital materials, such as f-electron systems.
@@
Figure 3: The comparison of the solution of the Anderson lattice model using ED and NQS, combining with the gGA quantum embedding method. The band spectrum and density of states (DOS) are displayed. The convergence criteria are 1e-3 for both methods. The NQS impurity solver reproduces the phase difference correctly.
@@ 3.  Benchmark Calculations of the ALM
In this Section we present gGA benchmark calculations on the single-orbital ALM on a Bethe lattice, see Eq. (5), within the limit of infinite coordination number, with a semicircular free density of states:
@@
From now on we set the half-bandwidth $D$ as the energy unit.
We evaluate the performance of our NQS-based impurity solver, which exhibits a rich interplay between Mott physics and the hybridisation between correlated and itinerant degrees of freedom, controlled by the on-site energy shift $\epsilon_{p}$ and the Hubbard interaction $U$.
To benchmark the accuracy of our NQS solver, we fixed $\epsilon_{p}=-1$ and considered two representative values of $U$: $U=0.5$ (metallic phase) and $U=3$ (Mott insulating phase), assessing the performance of our method in both regimes.
Our results are shown in Fig. 3 and Table. 1. In Fig. 3 is shown the gGA spectral function, see Eq. (9):
$\mathcal{A}(\epsilon,\omega)=-\frac{1}{\pi}\text{ImTr}\left[\mathcal{G}(\epsilon,\omega)\right]\,$, (9)
and the corresponding density of states:
@@
The Green’s functions are computed from the non-interacting Hamiltonian. At $U=0.5$, the spectral function clearly shows metallic behaviour with finite density at the Fermi level. On the other hand, at $U=3.0$, a gap appears, characteristic of a Mott insulator. By comparing results using an ED solver, it is confirmed that our NQS solver accurately captures the qualitative differences between these phases of the model.
Quantitative occupation number results are presented in Table. 1. Both impurity (d) and bath (p) orbitals exhibit fractional occupations consistent with metallic behaviour, showing minimal numerical differences (around $10^{-3}$), compared to ED. While in the insulating phase ($U=3$), the strong correlation fixes the electron in the $p$ orbital, making it almost fully filled. We observe that both ED and NQS methods have an almost fully occupied $p$ orbital, which explains the insulating gap in the spectrum function. The NQS solver shows slightly larger numerical differences in the impurity orbital occupancy in the insulating phase (as compared to ED), primarily due to stringent numerical accuracy requirements imposed by the gGA algorithm in strongly correlated regimes.
Table 1: The orbital occupation number computed at $U=0.5$ and $U=3.0$ by exact diagonalisation (ED) and neural quantum state (NQS) solver.
@@
Overall, our results showcase the effectiveness and suitability of the NQS solver within the gGA quantum embedding framework. Specifically, for gGA, the favourable computational scaling properties of NQS suggest potential for extending calculations to impurity models with larger orbital counts, potentially paving the way for applications to multi-orbital complex strongly correlated material systems beyond the reach of other methods.
III.1 Error control analysis.
@@
In this section, we present the error control metric E-tol and P-tol by comparing the convergence and time cost during independent ghost-gutzwiller quantum embedding loops using the NQS impurity solver. In this analysis, we adopted $U=0.5$ metal phase Anderson lattice model discussed in the last section as a running example. We set the bounds for error control metrics P-tol and E-tol as $\epsilon_{p}$ and $\epsilon_{t}$, which are chosen among ($10^{-3}$, $5\times 10^{-4}$, $10^{-4}$). All other settings, including the network structure, sizes, and the VMC optimisation parameters, are kept consistent in all tests. The analysis results are displayed in Fig.4. From this, we verified the joint importance of both E-tol and P-tol. We also observe considerable growth of the time cost when reducing the bound $\epsilon_{p}$ for P-tol.
Figure 4: The computational time cost and error analysis. (a)-(c): The residual during each self-consistent iteration of embedding loops using different error control parameters. (d) The time cost statistics during property sampling with different sampling accuracy.
First, we display the residual computed from the maximum absolute difference of the gGA parameters $\mathcal{R}_{i}$ and $\Lambda_{i}$ between the previous and current iterations. The results are arranged in subfigures (a)-(c) of Fig.4. In all cases, the residual decays rapidly in the first several iterations, followed by long and slow convergence to smaller values. The decrease of $\epsilon_{t}$ can assist in reducing the residual: for instance, for $\epsilon_{t}$$10^{-4}$ (dark red line), lower residual results, compared to other cases when fixing the $\epsilon_{p}$. Since this generally appears, we therefore conclude that the E-tol, which represents the wave function optimisation error, is important to the stability of the NQS VMC impurity solver. Moreover, comparing the three subfigures, we notice that the residual cannot be further reduced when $\epsilon_{p}$ is larger than $10^{-4}$, regardless of the choice of the bound for E-tol. For instance, in Fig. 4(a) with $\epsilon_{p}$ = $10^{-3}$, it appears that all samples cannot converge to a considerably smaller residual. This is caused by the instabilities of sampled properties, which lead to unstable quantum-embedding loops. While for $\epsilon_{p}$$10^{-4}$ (blue lines in Fig. 4 (c)), the convergence of all three samples with different $\epsilon_{t}$ is improved. This shows the importance of controlling the property sampling accuracy P-tol, in our method.
Nevertheless, a higher accuracy of the properties (smaller P-tol) comes with considerable computational cost. In Fig. 4 (d), we displayed the time cost analysis of the properties sampling. We keep the colour convention consistent with (a)-(c), where lighter ones indicate higher $\epsilon_{t}$ and darker ones indicate smaller $\epsilon_{t}$. We can clearly see the growth of the sampling time with the decrease of the bound of P-tol. We observe the time cost increases by over 100 times from $\epsilon_{p}$$10^{-3}$ to $10^{-4}$, while the increase is around 10 times to $\epsilon_{t}$$5\times 10^{-4}$. This aligns with the scaling of the VMC uncertainties $\epsilon$, where $\epsilon\propto O(1/\sqrt{M})$. In this case, a decreased error tolerance $\epsilon$ by an order of magnitude would increase $M$ by about 100 times. Comparably, the time cost for the wave function optimisation appears negligible, which only takes (1.63, 2.31, 5.70) seconds on average for $\epsilon_{t}$ among ($10^{-3}$, $5\times 10^{-4}$, $10^{-4}$). From this, we also observe that the growth of computational time with the decrease of the bound of E-tol is rather gentle, compared with that of the parameter P-tol.
@@
To conclude, the error analysis demonstrates that E-tol and P-tol are both important metrics for a numerically accurate solution. The dominant computational time cost is the properties sampling rather than the wave function optimisation. This finding calls for a better VMC sampling method to be innovated for quickly extracting the accurate physical observables from the NQS.
@@ 4.  Discussion and summary
Employing NQS as an impurity solver offers several advantages. Due to the universal approximation capability of neural networks, accurately approximating the ground state wave function is feasible with properly designed network architectures, significantly reducing approximation errors common in other solvers. Furthermore, the favourable computational scaling properties of variational Monte Carlo (VMC) enable straightforward extensions to multi-orbital models relevant for both theoretical investigation and realistic material simulation. Typically, the neural network evaluation scales as $O(N^{2})$, while the Monte Carlo sampling scales as $O(N)$. With proper parallelisation on modern computers, such scalability suggests significant potential for studying complex systems such as strongly interacting f-orbital materials and the extended active-space models. This provides valuable possibilities for advancing research on strongly correlated electron systems.
@@
Our study indicates that two remaining challenges need to be solved for further developing the most robust and general-purpose NQS impurity solver.
@@
  • Accelerated computation of embedding properties. Iterative updates in quantum embedding algorithms require precise computation of various physical observables. For example, the DMFT requires the Green’s Function as the matching target, while the DMET and the gGA rely on accurate evaluation of single-body and/or two-body reduced density matrices. In practice, the sampling accuracy of these properties strongly affects the convergence of the entire embedding algorithm. Insufficient sampling accuracy can result in violations of fundamental physical constraints, therefore potentially destabilising the iterative embedding cycle, as found in our analysis. Given the slow $O(1/\sqrt{M})$ convergence of the sampling accuracy, a naive sampling strategy requires very large computational resources - for instance, roughly $10^{8}$ samples to reach an accuracy of $10^{-3}$. Therefore, developing an enhanced or importance sampling method tailored specifically to NQS-VMC represents a useful research direction for enhancing computational efficiency and broadening the applicability of the NQS method.
@@
  • Incorporating spin Symmetries in NQS wave function. In strongly correlated electronic systems, it is often necessary to enforce specific spin symmetries (e.g., spin degeneracy, collinear, non-collinear spin configurations, etc). However, embedding these symmetries explicitly into NQS wave functions is nontrivial. Incorrect symmetries could lead to convergence issues and a physically incorrect solution. One practice, employed in this study, involves manually symmetrising computed properties post hoc. This approach, however, may introduce potential instability and convergence risk. Developing neural network architectures capable of inherently encoding spin symmetry should enhance the numerical stability and physical reliability of NQS impurity solvers.
@@
In summary, we have designed and benchmarked an NQS-based impurity solver with an error control system. When integrated with the ghost Gutzwiller Approximation on the Anderson Lattice Model, our solver accurately captures the key electronic properties of both its metallic and Mott insulating phases, yielding results in excellent agreement with exact diagonalisation. Importantly, our analysis of the computational budget identifies the primary bottleneck not as the wave function optimisation, but as the high-accuracy property sampling required by the embedding loop. This finding, therefore, establishes the development of efficient, high-precision sampling methods as the primary challenge to establishing NQS as a practical impurity solver within quantum embedding methods.

### v2

@@ II.3 The gGA Algorithmic Structure
In this work we apply the ghost-Gutzwiller approximation (gGA) to the single-orbital Anderson lattice model:
$\hat{H}$ $=\sum_{\langle i,j\rangle}\sum_{\sigma}\big(t_{ij}+\delta_{ij}\tau_{p}\big)\,\hat{p}^{\dagger}_{i\sigma}\hat{p}_{j\sigma}$
$\quad+V\sum_{i\sigma}\Big(\hat{p}^{\dagger}_{i\sigma}\hat{d}_{i\sigma}+\text{H.c.}\Big)$
$\quad+\mu\sum_{i}\hat{N}_{i}+\sum_{i}\hat{H}_{\mathrm{loc}}^{i},$ (5)
$\hat{H}_{\mathrm{loc}}^{i}$ $=\frac{U}{2}\left(\hat{n}_{di}-1\right)^{2}.$ (6)
Here $\{\hat{p}_{i\sigma}\}$ denote the itinerant (uncorrelated) degrees of freedom and $\{\hat{d}_{i\sigma}\}$ the interacting states, $\hat{n}_{di}=\sum_{\sigma}\hat{d}^{\dagger}_{i\sigma}\hat{d}_{i\sigma}$, and $\hat{N}_{i}=\hat{n}_{pi}+\hat{n}_{di}$ with $\hat{n}_{pi}=\sum_{\sigma}\hat{p}^{\dagger}_{i\sigma}\hat{p}_{i\sigma}$. The term $\hat{H}_{\mathrm{loc}}^{i}$ is the Hamiltonian of the $i$-th fragment.
The gGA is a variational framework based on the ansatz $|\Psi_{G}\rangle=\hat{\mathcal{P}}_{G}|\Psi_{0}\rangle$, where $|\Psi_{0}\rangle$ is a Slater determinant defined in an enlarged auxiliary Fock space and $\hat{\mathcal{P}}_{G}=\prod_{i}\hat{\mathcal{P}}_{i}$ is a product of local embedding maps into the physical Hilbert space [36, 21]. The variational energy is evaluated within the so-called Gutzwiller approximation [27], which becomes exact in the infinite-coordination-number limit.
As explained in Appendix, the gGA energy minimization problem can be cast in the form of a quantum-embedding (QE) framework [36, 21]. In particular, for single-band models such as the ALM, the resulting algorithmic structure requires computing iteratively the ground state of an EH, consisting of the local fragment Hamiltonian $\hat{H}_{\mathrm{loc}}^{i}$ coupled to a finite bath of $B$ auxiliary orbitals:
$\hat{H}^{i}_{\mathrm{emb}}$ $=\hat{H}_{\mathrm{loc}}^{i}$
$+\sum_{a=1}^{B}\sum_{\sigma}\Big([\mathcal{D}_{i}]_{a}\,\hat{d}_{i\sigma}^{\dagger}\hat{b}_{ia\sigma}+\text{H.c.}\Big)$
$+\sum_{a,b=1}^{B}\sum_{\sigma}[\Lambda_{i}^{c}]_{ab}\,\hat{b}_{ib\sigma}\hat{b}^{\dagger}_{ia\sigma}\,$, (7)
where $\{\hat{b}_{ia\sigma}\}$ are bath modes and $(\mathcal{D}_{i},\Lambda_{i}^{c})$ are updated self-consistently. The parameter $B$ controls the size of the variational space and, as shown in Refs. [36, 21, 41, 42], setting $B=3$ is generally sufficient to calculate ground-state properties to DMFT-level accuracy.
[…]
$d_{ab\sigma}$ $=\langle\Phi_{i}|\hat{b}_{ib\sigma}\hat{b}^{\dagger}_{ia\sigma}|\Phi_{i}\rangle$ (8)
$h_{b\sigma}$ $=\langle\Phi_{i}|\hat{d}^{\dagger}_{i\sigma}\hat{b}_{ib\sigma}|\Phi_{i}\rangle\,.$ (9)
This operation is the computational bottleneck of the gGA workflow, and motivates the development of scalable impurity solvers such as the NQS approach introduced here.
Symmetry Constraints
In Sec. III we report gGA calculations performed enforcing spin-rotational invariance at the level of the variational ansatz. As shown in previous work [43], this requirement can be formulated in the quantum-embedding representation by demanding that the embedding problem (fragment plus bath) is treated in a manner consistent with global SU(2) spin rotations, i.e., by selecting the singlet solution of the EH. In practice, rather than implementing an explicit restriction to the $S^{2}=0$ sector, we enforce this condition by adding a positive penalty term to the EH,
$\hat{H}^{i}_{\mathrm{emb}}\ \longrightarrow\ \hat{H}^{i}_{\mathrm{emb}}+\alpha\,\hat{\mathbf{S}}^{2},\qquad\alpha>0\,$, (10)
where $\hat{\mathbf{S}}$ is the total spin operator of the full embedding problem (including both fragment and bath). This shifts non-singlet states to higher energy and stabilizes the selection of the singlet ground state. In addition, to reduce statistical noise and improve numerical stability, we apply a post-processing symmetrization of the spin structure of the one-body density matrices that enter the gGA self-consistency update.
As shown in Refs. [36, 21], the symmetry condition that the gGA wavefunction is particle-number conserving (normal-state) is enforced by restricting the EH to the half-filled sector of the corresponding impurity++bath Hilbert space. Within our NQS impurity solver, this is enforced by restricting the Monte Carlo sampling to configurations with the corresponding total particle number (number-conserving updates).
@@
Figure 3: The comparison of the solution of the Anderson lattice model using ED and NQS, combining with the gGA QE method. The band spectrum and density of states (DOS) are displayed. The convergence criteria are 1e-3 for both methods. The NQS impurity solver reproduces the phase difference correctly.
@@ 3.  Benchmark Calculations of the ALM
In this Section we present gGA benchmark calculations on the single-orbital ALM, see Eq. (5), which exhibits a rich interplay between Mott physics and the hybridization between correlated and itinerant degrees of freedom, controlled by the on-site energy shift $\tau_{p}$ and the Hubbard interaction $U$.
The Bethe lattice
We focus on a Bethe lattice, within the limit of infinite coordination number, whose hopping-matrix eigenvalues have a semicircular free density of states:
@@
The spectral function of the ALM can be expressed as a function of the energy variable $\epsilon$, as follows:
$\mathcal{A}(\epsilon,\omega)=\lim_{\eta\rightarrow 0^{+}}-\frac{1}{\pi}\,\mathrm{Im}\,\mathrm{Tr}\!\left[\mathcal{G}(\epsilon,\omega+i\eta)\right]\,$, (12)
where $\mathcal{G}$ is the Green’s function, which can be calculated within the gGA formalism as discussed in the Appendix. In practice, we plot the broadened quantity obtained by evaluating the right-hand side at a small but finite $\eta$; in Fig. 3 we use $\eta=0.06$ (in units of the half-bandwidth). The corresponding density of states is obtained by averaging over $\epsilon$:
@@
We refer the reader to Appendix for the explicit expression of $\mathcal{G}$ in terms of the converged gGA variational parameters.
III.1 Benchmark calculations
Throughout this section we set the half-bandwidth $D$ as the energy unit, $\tau_{p}=-1$, and consider two physical regimes: $U=0.5$ (metallic phase) and $U=3$ (Mott insulating phase).
[…]
As a first benchmark, we compare the spectral functions obtained with the higher-accuracy NQS setting in Fig. 3. At $U=0.5$, the spectral function shows metallic behaviour with finite density at the Fermi level, while at $U=3.0$ a gap opens, characteristic of a Mott insulator. By comparison with ED, the NQS solver captures the qualitative differences between these two regimes.
We next quantify the agreement through the occupation numbers reported in Table 1. In the metallic regime ($U=0.5$), both impurity ($d$) and bath ($p$) occupations are reproduced accurately by the NQS solver, with differences on the order of $10^{-3}$ relative to ED. In the insulating regime ($U=3$), both ED and NQS yield an almost fully occupied $p$ orbital, consistent with the insulating gap. However, the impurity occupation $n_{d}$ is more sensitive to numerical accuracy in this regime: the default NQS setting shows a larger deviation/uncertainty compared to ED, while the higher-accuracy setting NQS(hac) yields improved agreement (Table 1). At a high level, this increased sensitivity in the Mott regime can be traced to the self-consistent structure of quantum-embedding updates, where small solver inaccuracies can propagate more strongly than in the metallic regime; a brief discussion of this mechanism is provided in Appendix.
Table 1: The orbital occupation number computed at $U=0.5$ and $U=3.0$ by exact diagonalization (ED) and neural quantum state (NQS) solver. NQS(hac) stands for higher-accuracy, in which we use stricter error control E-tol and P-tol during the solve.
@@ III.1 Benchmark calculations
Overall, our results showcase the effectiveness and suitability of the NQS solver within the gGA QE framework. Specifically, for gGA, the favorable computational scaling properties of NQS suggest potential for extending calculations to impurity models with larger orbital counts, potentially paving the way for applications to multi-orbital complex strongly correlated material systems beyond the reach of other methods.
III.2 Error control analysis.
@@
In this section, we present the error control metric E-tol and P-tol by comparing the convergence and time cost during independent gGA QE loops using the NQS impurity solver. In this analysis, we adopted $U=0.5$ metal phase Anderson lattice model discussed in the last section as a running example. We set the bounds for error control metrics P-tol and E-tol to be among ($10^{-3}$, $5\times 10^{-4}$, $10^{-4}$). All other settings, including the network structure, sizes, and the VMC optimization parameters, are kept consistent in all tests. The analysis results are displayed in Fig.4. From this, we verified the joint importance of both E-tol and P-tol. We also observe considerable growth of the time cost when reducing the P-tol.
[…]
@@
To conclude, the error analysis demonstrates that E-tol and P-tol are both important metrics for a numerically accurate solution. The dominant computational time cost is the properties sampling rather than the wave function optimization. This finding calls for a better VMC sampling method to be developed for quickly extracting the accurate physical observables from the NQS.
@@ 4.  Discussion and summary
Employing NQS as an impurity solver offers several advantages. Due to the universal approximation capability of neural networks, accurately approximating the ground state wave function is feasible with properly designed network architectures, significantly reducing approximation errors common in other solvers. Furthermore, the favourable computational scaling properties of variational Monte Carlo (VMC) enable straightforward extensions to multi-orbital models relevant for both theoretical investigation and realistic material simulation. Typically, the neural network evaluation scales as $O(N^{2})$, while the Monte Carlo sampling scales as $O(N)$ [16]. With proper parallelisation on modern computers, such scalability suggests significant potential for studying complex systems such as strongly interacting f-orbital materials and the extended active-space models. This provides valuable possibilities for advancing research on strongly correlated electron systems.
@@
Our study indicates that there are remaining challenges that need to be solved for further developing the most robust and general-purpose NQS impurity solver.
@@
  • Accelerated computation of embedding properties. Iterative updates in QE algorithms require precise computation of various physical observables. For example, the DMFT requires the Green’s Function as the matching target, while the DMET and the gGA rely on accurate evaluation of single-body and/or two-body reduced density matrices. In practice, the sampling accuracy of these properties strongly affects the convergence of the entire embedding algorithm. Insufficient sampling accuracy can result in violations of fundamental physical constraints, therefore potentially destabilising the iterative embedding cycle, as found in our analysis. Given the slow $O(1/\sqrt{M})$ convergence of the sampling accuracy, a naive sampling strategy requires very large computational resources - for instance, roughly $10^{8}$ samples to reach an accuracy of $10^{-3}$. Therefore, developing an enhanced or importance-sampling method tailored specifically to NQS-VMC represents a useful research direction for enhancing computational efficiency and broadening the applicability of the NQS method. In addition, it would also be useful to benchmark the stability and numerical accuracy requirements of different QE frameworks. For instance, workflows such as NQS-DMET and NQS-DMFT could have different requirements for sampling accuracy.
@@
  • Incorporating spin symmetries in NQS wave function. In strongly correlated electronic systems, it is often necessary to enforce specific spin symmetries (e.g., spin degeneracy, collinear, non-collinear spin configurations, etc). However, embedding these symmetries explicitly into NQS wave functions is nontrivial. Incorrect symmetries could lead to convergence issues and a physically incorrect solution. One practice, employed in this study, involves modifying the Hamiltonian with spin symmetry-related penalties and post-hoc symmetrisation. While robust in this study, it may also be beneficial to incorporate the symmetry inherent in the wave function ansatz. This may require symmetry-aware network structure design.
  • Exploring multiple t-U region and larger system. While the focus of the current work is to identify sampling issues and establish robust convergence criteria for NQS within a quantum-embedding framework, it is useful to apply it for larger systems and scan a wider range of interacting and hopping parameters.
@@
In summary, we have designed and benchmarked an NQS-based impurity solver with an error control system. When integrated with the ghost Gutzwiller Approximation on the Anderson Lattice Model, our solver accurately captures the key electronic properties of both its metallic and Mott insulating phases, yielding results in excellent agreement with exact diagonalization. Importantly, our analysis of the computational budget identifies the primary bottleneck not as the wave function optimization, but as the high-accuracy property sampling required by the embedding loop. This finding, therefore, establishes the development of efficient, high-precision sampling methods as the primary challenge to establishing NQS as a practical impurity solver within QE methods.